Recent Log Entries

Beam restore after RF worka

Log Number: 4130376Logbook(s): Entry Makers: carlinoBody: 

Sent upto 60 uA to Hall C. Beam upto 50 uA runs without problems. Above 55 uA limited by 8E02 BLM trips. Relative orbits through the machine match ZPOS from 08feb2023 23:22:58 with C beam.
Hall C @ 8E04 5.0. Hall A beam @ 11.6.
Relative orbits through the machine match ZPOS from 08feb2023 23:22:58 with A beam with minor orbital difference in 4R, but nominally the same throught he rest of the machine.
45 uA to Hall A without trips.
Lambertson Viewer was almost unusable. No sure if the camera is fried or what. No wires visible. Barely can see beam.

Current push

Log Number: 4130369Logbook(s): Entry Makers: dchriBody: 

Hall A/B/D (No Current)
Hall C (Current) ARC & EXT Orbit Locks ON. Energy locks on

(Fig.1) General setup for Hall C
(Fig.2) 40uA without trip
(Fig.2) 60uA immediate 8E02 trip
(Fig.2) 55uA RF trip
(Fig.3) 20uA RF trip
(Fig.3) 30uA without trip
(Fig.3) 40uA BLM trip
(Fig.4) 40uA without trip
(Fig.4) 45uA without trip
(Fig.4) 50uA BLM &/or RF trip
(Fig.5) 50uA without trip
(Fig.5) 55uA 8E02 trip
(Fig.6) 50uA RF trip after ~3.5min
(Fig.7) 40uA IonChamber trip
(Fig.8) 50uA 8E02 trip
(Fig.9) 50uA without trip
(Fig.9) 55uA RF trip

Hall B/C/D (No Current)
Hall A (Current) ARC & EXT Orbit Locks OFF.

iCalibrate: Hall C

Log Number: 4130365Logbook(s): Body: Calibration ParametersHall:CTarget: He-3 10cmPass: 5Note: Calibrated On:2023-Jan-13 22:57Calibrated By: mockFile:C-2023-01-13_2257_He-3_10cm_Pass-5.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)Compton3,830

Downtime Incident CLOSED: 2L26-4 repairs

Log Number: 4130362Logbook(s): Body: IDDescriptionPeriodCause11589Title:
2L26-4 repairs

Summary:
replace chassis to reduce tripsDuration:
3 hours 17 minutes

Time Down:
13-Feb-2023 09:07

Time Up:
13-Feb-2023 12:24

COMMENT ON Pump maintenance

Log Number: 4130361Logbook(s): Body: 

replaced tip seals.
2-13-2023
serial number - MY20340001
0.0 mtorr
pulled from 2L23 JAN 2023

Current push - Viewer checks

Log Number: 4130356Logbook(s): Entry Makers: dchriBody: 

Lambertson was not functioning properly...

(Fig.1) 8E02 viewer w/ Hall C
(Fig.2) 8E02 viewer w/ Hall A/C

iCalibrate: Hall A

Log Number: 4130351Logbook(s): Body: Calibration ParametersHall:ATarget: no targetPass: 4Note: Calibrated On:2023-Feb-08 01:15Calibrated By: kpriceFile:A-2023-02-08_0115_no_target_new_Pass-4.hcdIon ChamberExisting Setpoing (rads/hr)New Setpoint (rads/hr)Compton2,049</

Pages